answer the four question please

A built-in beam AB with bending stiffness El is loaded as shown and is further supported by a vertical bar at point B. The bar has a Youngs Modulus Ev, a length Lh and an area Ab. Bar a. What is the vertical reaction at A? b. What is the axial load in the bar? c. What is the vertical deflection at B?
Answer #1

The problem can be solved using force method cosndiering the reactio force on beamat B as redundant

Deflection of point B due to q on the cantilever beam AB = qL4/8EI

deflection of point B due to concentrated force P at B = PL3/3EI

Let the net displacement at B be \Delta

Force in bar = (EbAb/Lb)*\Delta

net displacement is given by

qL4/8EI - [(EbAb/Lb)*\Delta]L3/3EI = \Delta

qL4/8EI = \Delta [1+(EbAb/Lb)L3/3EI]

\Delta = (qL4/8EI )/[1+(EbAb/Lb)L3/3EI]

axial load in bar =(EbAb/Lb)*\Delta=(EbAb/Lb)* (qL4/8EI )/[1+(EbAb/Lb)L3/3EI]

vertical reaction at A=qL-axial lpoad in bar = qL-(EbAb/Lb)* (qL4/8EI )/[1+(EbAb/Lb)L3/3EI]

vertical deflection at B=\Delta=(qL4/8EI )/[1+(EbAb/Lb)L3/3EI]